# Comparing Specialised Small and General Large Language Models on Text Classification: 100 Labelled Samples to Achieve Break-Even Performance

### v2

@@ 4 Conclusion
The experiments presented in this paper used significant compute resources as they required multiple training and evaluation runs of multiple models (to deal with variance in results), as well as using large language models that requires a lot of computation even just for the inference. Overall, the experiments were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCO2eq/kWh (default value used as the actual efficiency of our HW instance was not measured). A cumulative of 4000 hours of computation was performed on hardware of type A100 PCIe 40GB (TDP of 250W). Total emissions are estimated to be 432 kgCO2eq of which 0 percents were directly offset. This does not include the compute used by the ChatGPT model behind API as we are not able to estimate these resources. These estimations were conducted using the MachineLearning Impact calculator presented in Lacoste et al. (2019). Whenever possible, we tried to reduce the compute resources used as much as possible. The most compute resources were used by the large language models – LLaMA-2, ChatGPT, Mistral-7B and Zephyr-7B. As the prompting and in-context learning with these models resulted in quite stable results, we decided to evaluate them only on a single setting (using 1 000 labelled training samples) and using a fraction of the whole test set (1 000 test samples). In addition, for the ChatGPT model, we evaluate only on 10 runs. Even in their reduced evaluation, these experiments used large fraction of the GPU hours. The most significant contributor was the instruction-tuning, especially with the medium sized models (Mistral/Zephyr), where we opted to reduce the number of steps and epochs used for the training.
Acknowledgements

### v3

@@ Appendix D Impact Statement: CO2 Emissions Related to Experiments
The experiments presented in this paper used significant compute resources as they required multiple training and evaluation runs of multiple models (to deal with variance in results), as well as using large language models that requires a lot of computation even just for the inference. Overall, the experiments were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCO2eq/kWh (default value used as the actual efficiency of our HW instance was not measured). A cumulative of 4000 hours of computation was performed on hardware of type A100 PCIe 40GB (TDP of 250W). Total emissions are estimated to be 432 kgCO2eq of which 0 percents were directly offset. This does not include the compute used by the GPT model behind API as we are not able to estimate these resources. These estimations were conducted using the MachineLearning Impact calculator presented in Lacoste et al. (2019).
Whenever possible, we tried to reduce the compute resources used as much as possible. The most compute resources were used by the large language models – LLaMA-2, LLaMA-3, GPT-4, Mistral-7B and Zephyr-7B. As the prompting and in-context learning with these models resulted in quite stable results, we decided to evaluate them only on a single setting (using 1 000 labelled training samples) and using a fraction of the whole test set (1 000 test samples). In addition, for the GPT model, we evaluate only on 10 runs. Even in their reduced evaluation, these experiments used large fraction of the GPU hours. The most significant contributor was the instruction-tuning, especially with the medium sized models (Mistral/Zephyr), where we opted to reduce the number of steps and epochs used for the training.